OpenStudy (anonymous):

What is - 2 3/8 expressed as an improper fraction?

OpenStudy (anonymous):

This should be placed in Math. Anyway, \(\Large -2\frac {3}{8}=-\frac{2\times 8+3}{8}=-\frac{19}{8}\)